### v1

# Why the Agent Made that Decision: Explaining Deep Reinforcement Learning with Vision Masks

Abstract
Due to the inherent lack of transparency in deep neural networks, it is challenging for deep reinforcement learning (DRL) agents to gain trust and acceptance from users, especially in safety-critical applications such as medical diagnosis and military operations. Existing methods for explaining an agent’s decision either require to retrain the agent using models that support explanation generation or rely on perturbation-based techniques to reveal the significance of different input features in the decision making process. However, retraining the agent may compromise its integrity and performance, while perturbation-based methods have limited performance and lack knowledge accumulation or learning capabilities. Moreover, since each perturbation is performed independently, the joint state of the perturbed inputs may not be physically meaningful. To address these challenges, we introduce VisionMask, a standalone explanation model trained end-to-end to identify the most critical regions in the agent’s visual input that can explain its actions. VisionMask is trained in a self-supervised manner without relying on human-generated labels. Importantly, its training does not alter the agent model, hence preserving the agent’s performance and integrity. We evaluate VisionMask on Super Mario Bros (SMB) and three Atari games. Compared to existing methods, VisionMask achieves a 14.9% higher insertion accuracy and a 30.08% higher F1-Score in reproducing original actions from the selected visual explanations. We also present examples illustrating how VisionMask can be used for counterfactual analysis.
1 Introduction
Original State LIME RISE Greydanus et al. SARFA VisionMask
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Table 1: Visual comparison of saliency maps generated by VisionMask, LIME(Ribeiro, Singh, and Guestrin 2016), RISE(Petsiuk, Das, and Saenko 2018), Greydanus et al. (2018), and SARFA(Puri et al. 2020) for the PPO agent(Schulman et al. 2017; Nguyen 2020) trained in the Super Mario Bros.(Kauten 2018) environment. Saliency level ranges from blue(low) to red(high). Row 1: The chosen action is to move left. The explanation is that Mario is attempting to avoid the Piranha Plants on top of the green pipes. VisionMask accurately identified the two Piranha Plants on the pipes. While RISE and Greydanus highlighted almost the entire image and LIME and SARFA highlighted areas not directly related to the explanation. Row 2:The chosen action is to jump right with acceleration. The explanation is that Mario intends to land on top of the Koopas to clear his path and get reward. Again VisionMask clearly highlighted the two Koopas whereas the other saliency maps failed to identify the correct region.
Deep Reinforcement Learning (DRL) is a powerful technology in machine intelligence, widely used for many applications, including ChatGPT (Brown et al. 2020; Achiam et al. 2023) and Alpha Go Zero (Silver et al. 2017). It enables effective end-to-end training of decision models that map an agent’s observations to actions, aiming to maximizing the long term expected rewards. However, understanding a DRL agent’s decision-making process is challenging, due to the inherent lack of explainability in the high-dimensional, non-linear structure of its underlying Deep Neural Network (DNN). The lack of transparency undermines human users’ trust and acceptance, particularly in safety-critical scenarios, such as medical diagnosis and military operations. To ensure wide acceptance and efficient use of this technology, it is crucial to provide accurate and human-interpretable explanations. These considerations motivated the recent interests in Explainable Reinforcement Learning (XRL).
Some existing works integrate explanation generation into the agent model and train them together. We refer to these as model-based approaches. For example, Interpretable DQN (i-DQN) (Annasamy and Sycara 2019) implements a deep Q-network with an attention mechanism, using the attention weights between state and action embeddings to represent their importance. Saliency-guided Q-networks (SGQNs) (Bertoin et al. 2022) calculate the importance of inputs as a derivative of the Q-function and train the critic model to both minimize the value prediction error and maximize the fidelity of the explanation. Some other works moved away from DNN and adopt special policy architectures that are inherently more interpretable (Liu et al. 2018; Verma et al. 2018; Bastani, Pu, and Solar-Lezama 2018). However, these models usually experience performance loss compared to the DNN-based policy model due to the limitations imposed by the special structure. Those model-based approaches cannot provide explanations for a trained DRL agent, unless they retrain the agent using their specific models, which is prohibitively expensive. And the retrain process may compromise the agent’s performance and integrity as it strives to maintain a balance between the optimality and explainability.
One of the most widely used techniques to provide explanations for machine learning models is the attribution-based method, which assigns credits to input features based on their contribution to AI agent’s decisions. This technique was originally developed to explain image classification models in computer vision. It generates saliency maps by perturbing the input image to reveal their contributions to the classification process without modifying the classifier itself (Dabkowski and Gal 2017; Petsiuk, Das, and Saenko 2018; Ribeiro, Singh, and Guestrin 2016). Inspired by this, several works have explored to use attribution-based technique to explain DRL models while treating the agent as a black-box. Greydanus et al. (2018) perturb the input features using Gaussian blur at fixed intervals and measure how the perturbation changes the action probability. The difference in action probability before and after perturbation are used to generate a saliency-map for the importance of each input feature. Similarly, Iyer et al. (2018) and Puri et al. (2020) compute the importance score based on the differences in Q values before and after the perturbation. The latter also improve the accuracy of the saliency map by normalizing the effects of perturbation between target and non-target actions.
The perturbation-based methods do not alter the agent model, nor do they impose any specific requirements on it. Hence they are applicable to any trained DRL agents. However, these techniques have some significant limitations. First, to manage complexity, perturbations are performed according to a set of predefined rules, such as using Gaussian blur or occluding with a constant value, which severely limits the search space they can explore. For DRL agents that map visual input directly to actions, due to the large input dimensions, there is no guarantee that the predefined perturbation can find the set of pixels that trigger the agent decision. Secondly, perturbation-based methods do not learn from past experiences, nor do they improve performance by comparing similarities and differences across different input states and actions. Thirdly, each input feature (e.g., input pixel) is perturbed independently, with no guarantee that the joint state of the perturbed pixels represents meaningful objects. As shown in Table 1, the saliency maps generated by the perturbation-based method either consider all pixels equally important or identify regions in the image that do not contain meaning objects.
To address the aforementioned challenges, we propose VisionMask, a model trained end-to-end to generate attribution-based explanations for vision-based DRL without perturbation or altering the agent. VisionMask ouputs a saliency map that highlights the critical regions in the agent’s visual input responsible for triggering the agent’s action. The saliency map reveals the importance of each pixel in the decision making process, and will also be referred to as importance map or importance mask in the rest of the paper. VisionMask is trained using self-supervised contrastive learning. No human generated label is required. The training of VisionMask does not alter the agent model, thereby preserving the agent’s performance and integrity. The (fixed) policy model of the DRL agent is integrated into the training pipeline to generate the loss, and involved in the backpropagation to pass the gradient to the mask generation model. Therefore, each VisionMask model is trained to explain a specific DRL agent. The training process is agent agnostic, meaning that it works with any agent with a policy model that maps visual input to actions. VisionMasks’s learnable capability allows it to accumulate knowledge from previous experiences to improve future explanations. During testing time, VisionMask does not need to access agent’s internal state to generate the importance mask. The mask is generated by a forward pass of the VisionMask model.
As shown in Table 1, VisionMask accurately identifies the critical regions in the visual input that explains the agent’s action. Beyond providing precise explanations directly from the visual input, VisionMask also facilitates counterfactual analysis. By selectively activating or deactivating different critical regions, meaningful counterfactual scenarios can be generated and assessed. This is similar to the perturbation-based approach, however, at a much higher level.
The contribution of our work can be summarized as the following:
  • We present VisionMask, a novel attribution method for explainable visual reinforcement learning that generates action-specific importance masks. Generated using a single forward pass, the mask does not only highlight the regions in the visual input that explain the agent’s behavior but also facilitates counterfactual analysis.
  • VisionMask is model-agnostic in that it works with any vision-based DRL agent without modifying the agent model.
  • VisionMask is trained using self-supervised contrastive learning. No human intervention is needed for data labeling.
  • The fidelity of the explanation is measured by the agreement between the agent’s actions generated from the original visual input versus from the masked visual input. We quantify this agreement using accuracy, precision, recall and F1-score. In average, VisionMask outperforms other attribution-based explanation models by 30.08% on F1-score and 14.9% on insertion accuracy..
2 Background and related work
The nested structure and non-linear operation of DNN make it challenging for humans to understand how the outputs are derived from inputs. Some existing works address this challenge by explaining model outputs through input attributions (Dabkowski and Gal 2017; Petsiuk, Das, and Saenko 2018; Ribeiro, Singh, and Guestrin 2016; Fong and Vedaldi 2017). For example, Randomized Input Sampling for Explanation of Black-box Models (RISE) (Petsiuk, Das, and Saenko 2018) is a perturbation-based approach that explains neural image classifiers by applying randomly generated masks to the input image and assessing their impact. Local Interpretable Model-agnostic Explanations (LIME)(Ribeiro, Singh, and Guestrin 2016) tries to explain local instances by approximating them within a nearby vicinity using a linear decision model, where the explainable elements are superpixels (i.e., small image regions of similar pixels).
Built on top of the DNNs, the lack of transparency of the DRL agents (Samek, Wiegand, and Müller 2017; Ras et al. 2022; Angelov and Soares 2020) undermines humans trust and hinders their adoption. Several studies have addressed this challenging problems. According to (Qing et al. 2022), existing efforts can be categorized into four main approaches: model explaining, reward explaining, task explaining, and state explaining.
Model explaining relies on inherently interpretable model architectures or auxiliary reasoning mechanisms to generate explanations (Topin et al. 2021; Jiang and Luo 2019; Landajuela et al. 2021). However, these self-explanatory or symbolic models often suffer from decreased performance compared to state-of-the-art neural network based RL policies and may lack the representational power needed to learn more complex policies. Reward explanation typically involves using explicitly designed explainable reward functions to generate explanations (Juozapaitis et al. 2019; Ashwood, Jha, and Pillow 2022). Task explanation considers a policy as a composition of sub-task and action sequences, explaining the behavior in terms of relationships among sub-tasks (Shu, Xiong, and Socher 2017; Sodhani, Zhang, and Pineau 2021). These approaches often assume the existence of a reward decomposition scheme or a predefined sub-task partition mechanism, which may not hold true in all RL environments.
State explanations are based on agent observations from the environment. This approach determines the significance of explainable elements within a state observation in relation to action selection or reward determination. An explainable element in the state could be a small region in the visual input or semantic features in the environment state. The proposed VisionMask falls into this category.
State explaining methods can further be divided into three categories: attention-based, Shapley value-based and perturbation-based mechanisms, which are detailed below. Attention is a common approach used for explainable RL in several existing works (Peng, Riedl, and Ammanabrolu 2022; Xu et al. 2020; Annasamy and Sycara 2019). However, similar to i-DQN (Annasamy and Sycara 2019), these methods cannot explain a given DRL model because their attention model must be trained alongside the agent model. Shapley value (Shapley et al. 1953), a concept from game theory, has also been introduced into XRL. SVERL (Beechey, Smith, and Şimşek 2023) provides a theoretical analysis on how Shapley values can be applied to explain value and policy network in DRL. However, this approach is still in its early stages and assumes prior knowledge of the transition model of the environment, which is not realistic for realistic applications.
Perturbation-based methods (Greydanus et al. 2018; Iyer et al. 2018; Puri et al. 2020) compute saliency maps of the input features by comparing the action probability or value function before and after perturbation. Greydanus et al. (2018) perturbed the state with Gaussian blur at fixed intervals, while Iyer et al. (2018) identified objects and perturbed them with a constant gray value. Puri et al. (2020) improved on Greydanus et al. (2018) by removing the effect of perturbations from other irrelevant actions. As discussed in section 1, the performance of these techniques is constrained by predefined perturbation rules and the lack of knowledge accumulation. Furthermore, there is no guarantee that the perturbed input remains physically meaningful.
3 VisionMask
This section presents VisionMask, an innovative agent-agnostic framework for XRL. The primary goal is to produce action-wise importance-mask that attribute explainable elements in the current state to each possible action. The definition of explainable elements may vary across different applications. However, in this paper, we focus on identifying explanatory pixel regions in the agent’s visual input.
Figure 1: Architecture of VisionMask. In this example $a=1$ so $m_{1}$ is the active mask $m_{a}$.
3.1 Problem Formulation
We approach the task of VisionMask as an imitation learning problem based on Behavioral Cloning (BC) (Bain and Sammut 1995; Ross, Gordon, and Bagnell 2011). Formally, we define the environment as a Markov Decision Process (MDP) $\{\mathcal{S},\mathcal{A},P,R,\gamma\}$, where $\mathcal{S}$ represents the state space; $\mathcal{A}$ denotes the action space; the state transition function $P:\mathcal{S}\times\mathcal{A}\to\Delta(\mathcal{S})$ depicts the transition between states based on actions, where $\Delta(\mathcal{S})$ represents the set of probability distributions over states; the reward function $r:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$ provides the immediate reward for state-action pairs; and $\gamma\in[0,1]$ and $\pi:\mathcal{S}\to\Delta(\mathcal{A})$ represent the discount factor and policy. Return $G$ is defined as $G=\sum_{k=0}^{\infty}\gamma^{k}R_{k+1}$, and the expected cumulative reward of a policy $\pi$ is $\mathbb{E}_{\pi}[G]=\mathbb{E}_{\pi}[\sum_{k=0}^{\infty}\gamma^{k}R_{k+1}]$, where the expectation is taken with respect to the initial state distribution, transition probabilities, and action probabilities determined by $\pi$.
VisionMask operates on a given pre-trained expert policy $\pi_{E}$ such that
$\pi_{E}\approx\pi^{*}=\operatorname*{arg\,max}_{\pi}\mathbb{E}_{\pi}[G],$
where $\pi^{*}$ is the optimal policy. We can obtain a dataset of expert demonstrations $\mathcal{D}_{E}=\{(s_{i},\pi_{E}(s_{i}))\}_{i=1}^{N}$ consisting of $N$ state-action pairs, from trajectories sampled while executing $\pi_{E}$ in the environment. Our goal is to learn an explainer $f_{\theta^{*}}$ that minimizes the loss
$\theta^{*}=\operatorname*{arg\,min}_{\theta}\sum_{(s,a)\in\mathcal{D}_{E}}L(a,s,\theta),$
where $L$ is the training loss function to be discussed in section 3.3. The explainer function $f_{\theta}:\mathcal{S}\to[0,1]^{K\times d_{s}}$ (with $d_{s}$ representing the state dimension and $K$ the number of candidate actions) provides the attributions of each action to each feature in the state $s$. In the case of visual DRL, $s$ is the visual input of the agent and $d_{s}$ is the number of pixels in $s$, i.e., $d_{s}=W\times{H}$, with $W$ and $H$ representing the width and height of the visual input. The value of each element in the attribution is bounded within the range $[0,1]$, with the $(i,j)$th element indicates the $j$th pixel’s attribution to the $i$th action. This attribution will then be overlaid onto the visual input, masking unimportant pixels while retaining important ones. The overlaid visual input will be analyzed by the policy model again. We will update parameters $\theta$ of explainer $f$ so that the overlaid visual input always produces the same action as the original one. The expert policy $\pi_{E}$ remains frozen during the training process of $f_{\theta}$, meaning its parameters are fixed.
3.2 Architecture
As shown in Fig. 1, the expert dataset $\mathcal{D}_{E}$ is collected using the expert policy $\pi_{E}$. From this dataset $\mathcal{D}_{E}$, state-action pairs $(s_{i},a_{i})$ are sampled. Subsequently, the explainer $f_{\theta}$ takes the state $s_{i}$ as input and produces a set of action-wise masks $M=\{m_{0},m_{1},....,m_{K-1}\}$, where $K$ denotes the number of candidate actions and $m_{i}\in{[0,1]^{W\times{H}}}$. A mask value $m_{i}[x,y]$ near 1 indicates that the pixel $(x,y)$ has significant contributions to the corresponding action $i$, whereas a value near 0 denotes irrelavance. This process shares similarities with image segmentation, where each pixel is attributed to an object or background. Hence, we structure the explainer $f_{\theta}$ akin to the widely used image segmentation model, DeepLabV3 (Chen et al. 2017), but retrain it with the expert demonstrations $\pi_{E}$. To make sure that the output masks are bounded to the range $[0,1]$, a sigmoid function is applied at the output of $f_{\theta}$
Following (Stalder et al. 2022), we partition the masks in $M$ into two sets: the active set and the inactive set. Unlike (Stalder et al. 2022), our active set has only one mask, $m_{a}$ where $a$ is the action selected by expert policy $\pi_{E}$ll the other masks, $m_{i}$, where $i\neq{a}$, belong to the inactive set. Each mask $m_{i}$ contains the attributions for the current agent state that explain the corresponding action $i$. The explainer will be trained to produce masks for pixels that contribute the most to the policy’s action selection.
We aggregate the masks in the inactive set into a non-target mask $n$ by taking the maximum value pixel-wise. We also calculate a complement mask $\tilde{m_{a}}=\mathbf{1}-m_{a}$ representing the less important regions for the selected action $a$. Then the masks $m_{a}$ and $\tilde{m_{a}}$ are overlaid onto the original state $s$ to generate two masked states $s_{m}$ and $s_{\tilde{m}}$ using the following overlay function:
$s_{m}=s\odot m_{a}+r\odot(\tilde{m_{a}}),s_{\tilde{m}}=s\odot\tilde{m_{a}}+r\odot(m_{a})$
where $\odot$ is Hadamard Product and $r$ is a reference value. The final step is to feed these masked states $s_{m}$ and $s_{\tilde{m}}$ through the expert policy $\pi_{E}$ to generates action distributions $\mathbf{p}$ and $\mathbf{\tilde{p}}$, respectively:
$\mathbf{p}=\pi_{E}(\mathbf{s}_{\mathbf{m}}),\mathbf{\tilde{p}}=\pi_{E}(\mathbf{s}_{\mathbf{\tilde{m}}}).$
The model $f_{\theta}$ is optimized so that $\mathbf{p}$ matches the action distribution chosen for the unmasked state $s$, while $\mathbf{\tilde{p}}$ does not.
It is worth mentioning that the agent may be sensitive to the reference value in the overlay function. Numerous options exist for the reference value, such as setting the pixel to zero, assigning a constant value, blurring the pixel, or cropping it. In this study, we opt for a constant value as the reference value, based on the sensitivity of the agent to zero and blurring, as indicated by experimental results. We use different reference values for different environments (See Appendix 4).
3.3 Training Loss
To train the explainer to capture the most important regions in the state, we carefully designed the training loss function $\mathcal{L}$ as follows:
$\mathcal{L}(\mathbf{s},\mathbf{a},\theta,\mathbf{n})$ $=\mathcal{L}_{bc}(\mathbf{s},\mathbf{a})+\lambda_{e}\mathcal{L}_{e}(\mathbf{s},\mathbf{a})+\lambda_{avg}\mathcal{L}_{avg}(\mathbf{s},\mathbf{a},\mathbf{n})$
$+\lambda_{\text{smooth}}\mathcal{L}_{\text{smooth}}(\mathbf{s},\mathbf{a},\mathbf{n})+\lambda_{L2}\lVert\theta\rVert^{2}_{2},$
where $\mathcal{L}_{bc}(\mathbf{s},\mathbf{a})$ is Behavior Cloning loss, $\mathcal{L}_{e}(\mathbf{s},\mathbf{a})$ is entropy loss, $\mathcal{L}_{avg}(\mathbf{s},\mathbf{a},\mathbf{n})$ is the mask average loss, $\mathcal{L}_{\text{smooth}}(\mathbf{s},\mathbf{a},\mathbf{n})$ is the smoothness loss, $\lVert\theta\rVert^{2}_{2}$ is $L_{2}$ regularization and $\lambda_{e},\lambda_{avg},\lambda_{\text{smooth}},\lambda_{L2}$ are regularization hyper-parameters.
Behavioral Cloning loss $\mathcal{L}_{bc}$. Our goal is to train the explainer to select regions in the visual input that that are sufficient and necessary for the agent to choose action $\mathbf{a}$. Hence, we design the behavior cloning loss as cross entropy loss between the action distribution $\mathbf{p}$ predicted for the overlaid state $s_{m}$ and the target $\mathbf{a}$ chosen for the original state $s$:
$\mathcal{L}_{bc}(\mathbf{s},\mathbf{a})=-\frac{1}{K}\sum_{k=1}^{K}\llbracket k=\mathbf{a}\rrbracket\log(\mathbf{p}[k]).$
Here $\llbracket k=\mathbf{a}\rrbracket$ denotes the indicator function which returns 1 if $k$ is the same as label $\mathbf{a}$, and 0 otherwise.
Negative Entropy loss $\mathcal{L}_{e}$. To ensure that the visual input regions selected by $m_{a}$ is sufficient and necessary for the agent to make decisions, we also need to make sure that the unselected region, i.e., $s_{\tilde{m_{a}}}$, does not provide useful information for action selection, hence the action distribution $\mathbf{\tilde{p}}$ should follow a uniform distribution. Motivated by this rationale, we define negative entropy loss regarding $\tilde{\mathbf{m}}$ as the following:
$\mathcal{L}_{e}(\mathbf{s},\mathbf{a})=\frac{1}{K}\sum^{K}_{k=1}\tilde{\mathbf{p}}[k]\log\tilde{\mathbf{p}}[k].$
Mask average loss $\mathcal{L}_{avg}$. A low effort way to minimize $\mathcal{L}_{bc}$ and $\mathcal{L}_{e}$ is to include all pixels in the importance mask, $m_{a}$, and no pixel in the complement mask, $\tilde{m_{a}}$, which obviously is not a valid solution. We need to ensure that each importance mask only consists a small number of crucial pixels. Thus, we define $\mathcal{L}_{avg}$ as follows:
$\mathcal{L}_{avg}(\mathbf{s},\mathbf{a},\mathbf{n})=\frac{1}{W\times H}\sum_{i,j}\mathbf{m}[i,j]+\frac{1}{W\times H}\sum_{i,j}\mathbf{n}[i,j],$
where, $\mathbf{n}$ represents the non-target mask aggregated from the inactive set. By minimizing the mask average loss, we limit the number of pixels selected by the masks.
Smoothness loss $\mathcal{L}_{\text{smooth}}$. Finally, we minimize the Variation between adjacent pixels to generate smoother masks:
$\mathcal{TV}(x)$ $=\frac{1}{W\times H}\sum_{i,j}(|x[i,j]-x[i+1,j]|$
$+|x[i,j]-x[i,j+1]|)$
$\mathcal{L}_{\text{smooth}}(\mathbf{s},\mathbf{a},\mathbf{n})=\mathcal{TV}(m)+\mathcal{TV}(n)$
The overall architecture of VisionMask is given in Fig. 1. The loss will backpropagate through the frozen expert model $\pi_{E}$ to train the explainer model $f_{\theta}$.
4 Experiments
Env Metric VisionMask LIME RISE Greydanus SARFA
Mario Acc. $\mathbf{0.9076}\pm\scriptstyle{0.0124}$ $0.7868\pm\scriptstyle{0.0027}$ $0.8053\pm\scriptstyle{0.0118}$ 0.1692 0.1692
Precision $\mathbf{0.7109}\pm\scriptstyle{0.0189}$ $0.2448\pm\scriptstyle{0.0200}$ $0.3364\pm\scriptstyle{0.0046}$ 0.6154 0.6154
Recall $\mathbf{0.7005}\pm\scriptstyle{0.0020}$ $0.2467\pm\scriptstyle{0.0042}$ $0.3245\pm\scriptstyle{0.0069}$ 0.1752 0.1800
F1-score $\mathbf{0.6995}\pm\scriptstyle{0.0058}$ $0.2277\pm\scriptstyle{0.0056}$ $0.3046\pm\scriptstyle{0.0082}$ 0.2764 0.2753
Enduro Acc. $\mathbf{0.9061}\pm\scriptstyle{0.0077}$ $0.8811\pm\scriptstyle{0.0087}$ $0.9033\pm\scriptstyle{0.0047}$ 0.2744 0.2790
Precision $0.6058\pm\scriptstyle{0.0084}$ $0.4167\pm\scriptstyle{0.0062}$ $0.5363\pm\scriptstyle{0.0054}$ $\mathbf{0.6449}$ 0.6383
Recall $\mathbf{0.5535}\pm\scriptstyle{0.0049}$ $0.4551\pm\scriptstyle{0.0046}$ $0.5245\pm\scriptstyle{0.0072}$ 0.0839 0.0889
F1-score $\mathbf{0.5060}\pm\scriptstyle{0.0082}$ $0.3809\pm\scriptstyle{0.0018}$ $0.4738\pm\scriptstyle{0.0032}$ 0.1283 0.1418
Seaquest Acc. $\mathbf{0.9531}\pm\scriptstyle{0.0030}$ $0.9094\pm\scriptstyle{0.0079}$ $0.9201\pm\scriptstyle{0.0051}$ 0.4463 0.4476
Precision $0.5809\pm\scriptstyle{0.0022}$ $0.2647\pm\scriptstyle{0.0066}$ $0.4631\pm\scriptstyle{0.004}$ 0.8083 $\mathbf{0.8371}$
Recall $\mathbf{0.4880}\pm\scriptstyle{0.0013}$ $0.2423\pm\scriptstyle{0.0082}$ $0.2830\pm\scriptstyle{0.0047}$ 0.0945 0.0983
F1-score $\mathbf{0.4714}\pm\scriptstyle{0.0050}$ $0.2073\pm\scriptstyle{0.01443}$ $0.2733\pm\scriptstyle{0.0110}$ 0.1640 0.1701
MsPacman Acc. $\mathbf{0.9768}\pm\scriptstyle{0.0177}$ $0.8451\pm\scriptstyle{0.0294}$ $0.9264\pm\scriptstyle{0.0220}$ 0.4616 0.4235
Precision $0.7671\pm\scriptstyle{0.0160}$ $0.1878\pm\scriptstyle{0.0298}$ $0.5590\pm\scriptstyle{0.0097}$ 0.8255 $\mathbf{0.9102}$
Recall $\mathbf{0.8658}\pm\scriptstyle{0.0039}$ $0.1726\pm\scriptstyle{0.0036}$ $0.5587\pm\scriptstyle{0.0035}$ 0.2189 0.1913
F1-score $\mathbf{0.7699}\pm\scriptstyle{0.0096}$ $0.1597\pm\scriptstyle{0.0086}$ $0.5122\pm\scriptstyle{0.0078}$ 0.3307 0.3067
Table 2: Fidelity scores on four environments and confusion matrices in the Mario environment. We run each experiment 3 times with different random seeds and report the mean and standard deviation.
We evaluated VisionMask across four different environments and compared its performance against four baseline methods.
Environment Selection. Four agent-centered RL environments were selected to showcase the effectiveness of VisionMask: Super Mario Bros (SMB)(Kauten 2018), and the Atari games Seaquest, Enduro, and MsPacman. We select these environments following existing works (Puri et al. 2020; Greydanus et al. 2018).
Baseline Selection. We compared VisionMask’s performance against four perturbation-based explainer, LIME (Ribeiro, Singh, and Guestrin 2016) RISE (Petsiuk, Das, and Saenko 2018), Greydanus et al. (2018) and SARFA(Puri et al. 2020). For LIME, we used the original public implementation with the default configuration available from (Ribeiro 2017). For RISE, we adopt the public implementation from torchray (Fong, Patrick, and Vedaldi 2019). For Greydanus et al. (2018) and SARFA, we used the respective public implementation provided by the authors. To ensure a fair comparison, we did not include any explainers that require retraining the agent model.
Expert Policy. For reproducibility, we did not train the expert policies ourselves; instead, we used publicly available implementations. In the SMB environment, We apply the trained PPO (Schulman et al. 2017) policy from (Nguyen 2020). In the Atari environments, we used three trained PPO policies from Stable-Baselines3 (Raffin et al. 2021) .
4.1 Assessment of Explanation Fidelity
We quantify fidelity using four metrics: accuracy, precision, recall, and F1-score. Fidelity measures the agreement between the agent’s action selected for the original visual input and the action selected for the masked visual input. In this context, we define the accuracy as the following:
$\text{Accuracy}=\frac{1}{|\mathcal{D}_{\text{test}}|}\sum_{i=1}^{|\mathcal{D}_{\text{test}}|}\llbracket\pi_{E}(s_{i})=\pi_{E}(s_{i}^{\prime})\rrbracket,$
[…]
Values of these fidelity metrics are shown in Table 2. VisionMask consistently scores higher than LIME and RISE. Greydanus and SARFA have a higher precision than VisionMask, however, a much lower recall and F1-score. This is probably because they included almost all pixels from the visual input in the importance map. They also have a lower accuracy compared to VisionMask.
Refer to caption Refer to caption
Original state VisionMask
Refer to caption Refer to caption
Insertion curve Deletion curve
Figure 2: Examples of insertion and deletion AUC curve
Env Method 0.25 0.5 1.0
$\downarrow$ del $\uparrow$ ins $\downarrow$ del $\uparrow$ ins $\downarrow$ del $\uparrow$ ins
Mario VisionMask 0.2044 ±0.1891 0.4156 ±0.2937 0.1813 ±0.1730 0.4142 ±0.2788 0.1714 ±0.1577 0.4216 ±0.2622
LIME 0.3447 ±0.2664 0.2314 ±0.2189 0.3231 ±0.2568 0.2446 ±0.2232 0.2693 ±0.1958 0.3024 ±0.2107
RISE 0.2750 ±0.2510 0.2940 ±0.2624 0.2331 ±0.2326 0.3317 ±0.2761 0.1892 ±0.1897 0.3796 ±0.2678
Greydanus 0.5708 ±0.2109 0.1963 ±0.1658 0.5671 ±0.2146 0.1952 ±0.1650 0.5634 ±0.2169 0.2018 ±0.1691
SARFA 0.5579 ±0.2223 0.2071 ±0.1756 0.5565 ±0.2235 0.2117 ±0.18 0.5556 ±0.2246 0.2143 ±0.1866
Enduro VisionMask 0.3693 ±0.2854 0.3952 ±0.3004 0.3589 ±0.2911 0.3979 ±0.2995 0.3491 ±0.2948 0.4021 ±0.2941
LIME 0.3780 ±0.2886 0.3636 ±0.3025 0.3723 ±0.2914 0.3680 ±0.3002 0.3629 ±0.2937 0.3799 ±0.2883
RISE 0.4081 ±0.2818 0.3585 ±0.3066 0.3970 ±0.2861 0.3714 ±0.2938 0.3759 ±0.2912 0.3906 ±0.2785
Greydanus 0.3762 ±0.2988 0.3312 ±0.3194 0.3626 ±0.3018 0.3317 ±0.3192 0.3457 ±0.3075 0.3329 ±0.3188
SARFA 0.3472 ±0.3060 0.3327 ±0.3188 0.3378 ±0.3114 0.3336 ±0.3186 0.3331 ±0.3157 0.3340 ±0.3186
Seaquest VisionMask 0.0721 ±0.0291 0.3256 ±0.2596 0.0684 ±0.0289 0.3331 ±0.2650 0.0652 ±0.0268 0.3430 ±0.2635
LIME 0.1982 ±0.1871 0.1397 ±0.1662 0.1909 ±0.1865 0.1425 ±0.1672 0.1434 ±0.1214 0.2141 ±0.1812
RISE 0.1198 ±0.0797 0.2422 ±0.2560 0.0987 ±0.0658 0.2638 ±0.2654 0.0809 ±0.0505 0.2896 ±0.2660
Greydanus 0.2939 ±0.2381 0.0658 ±0.0345 0.2165 ±0.1834 0.0708 ±0.0542 0.1439 ±0.0980 0.0788 ±0.0825
SARFA 0.1133 ±0.0732 0.0819 ±0.0978 0.0888 ±0.0418 0.0851 ±0.1104 0.0764 ±0.0307 0.0869 ±0.1177
MsPacman VisionMask 0.2251 ±0.1051 0.6350 ±0.3056 0.2037 ±0.1255 0.6231 ±0.3337 0.1956 ±0.1217 0.6278 ±0.2778
LIME 0.4831 ±0.2700 0.1901 ±0.1708 0.4348 ±0.2455 0.2346 ±0.1875 0.3265 ±0.1801 0.3771 ±0.1890
RISE 0.4270 ±0.3140 0.2299 ±0.1220 0.3512 ±0.2995 0.2937 ±0.1325 0.2621 ±0.2273 0.4577 ±0.1394
Greydanus 0.5669 ±0.2476 0.1285 ±0.0646 0.4903 ±0.2518 0.1336 ±0.0843 0.3905 ±0.2281 0.1417 ±0.1029
SARFA 0.3511 ±0.2528 0.1463 ±0.1168 0.2524 ±0.1573 0.1531 ±0.1328 0.1880 ±0.0873 0.1561 ±0.1425
Table 3: Insertion and deletion metrics for fractions 0.25, 0.5, and 1.0. The symbol $\uparrow$ (resp. $\downarrow$) indicates that higher (resp. lower) values are better. We report the mean $\pm$ standard deviation over testing samples.
4.2 Assessment of Effectiveness
An effective explainer should precisely identify the exact input regions that contribute to the expert decision, without including irrelevant areas. To quantify the effectiveness of explanations, We adopt the insertion and deletion metrics from RISE (Petsiuk, Das, and Saenko 2018). The rationale behind these metrics is as follows: If certain pixels in the visual state are crucial to the action (i.e., the true cause of the action), deleting them from the state or inserting them into the background should result in a corresponding decrease or increase in the probability of the specific action. By plotting the relationship between the probability of the target action and the number of pixels added or removed, we obtain an insertion curve and a deletion curve. The Area Under the Curve (AUC) of those plots represents the effectiveness of the explanation. A higher AUC for the insertion curve and a lower AUC for the deletion curve stand for more effective explanation. The order of insertion and deletion affects the result. We sort the pixels based on their values in the importance map and insert or delete them in descending order of their importance. Figure 2 shows an example of insertion and deletion curves with corresponding AUC values displayed at the top right corner. See Appendix D for the pseudocode of each metric.
It is worth noting that the effectiveness varies as different number of pixels are inserted or deleted from the original input. Instead of showing the actual curve for each explainer under different games, we report the AUC under 25%, 50%, and 100% level of insertion and deletion. The results are given in Table 3. VisionMask outperforms all other methods across all environments, with an average insertion score that is 14.9% higher and a deletion score that is 8.8% lower compared to the other methods.
4.3 Counterfactual Explanations
Original state VisionMask Counterfactual
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Move left Jump right
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Left Left + acceleration
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Down Fire Down right
Figure 3: Counterfactual analysis with VisionMask.
By removing the most important regions identified by VisionMask from the visual input, we can create scenarios for counterfacutal analysis. Table 3 provides three examples of this approach. In the first example, VisionMask correctly identified the Piranha Plants as the reason the agent chose to move left. After removing the Piranha Plant by replacing the highlighted pixels with reference value and run the modified image by the agent again, we observed that the agent’s action changed to jumping right, as the threat from the Piranha Plant no longer existed. In the second example, VisionMask correctly identified the edge of the road and a car in front as the reason why the agent chose to move left (with current speed). After we removed the car in front the agent chose to move left with acceleration. In the third example, VisionMask identified the shark in the bottom left as the cause of the downward and fire action. When the corresponding shark was removed, the agentno longer chose to fire. Appendix I, J and K contain more examples. We believe such explanations can further help users to understand the RL agent’s behavior.
5 Conclusion
We presented VisionMask, an agent-agnostic DRL explanation model trained in self-supervised contrastive learning. VisionMask generates explanations with higher fidelity and better effectiveness compared to existing attribute-based methods. It is our future plan to extend this approach to multi-modality input and couple the visual explanation generated by the VisionMake with other information such as agent’s long-term goals and future rewards.
Acknowledgements
This research is partially supported by the Air Force Office of Scientific Research (AFOSR), under contract FA9550-24-1-0078. The paper was received and approved for public release by AFRL on May 28th 2024, case number AFRL-2024-2908 . Any Opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of AFRLor its contractors.
Appendix A Technical Appendix
Appendix B Meanings of actions across four environments
In Table 4, we present the list of actions and their corresponding meanings for the four environments evaluated in the experiment section.
Value Meaning Value Meaning
0 Stay in place 0 No operation
1 Move right 1 Accelerate
2 Jump right 2 Move right
3 Move right with acceleration 3 Move left
4 Jump right with acceleration 4 Decelerate
5 Jump 5 Move right and decelerate
6 Move left 6 Move left and decelerate
7 Accelerate and move right
8 Accelerate and move left
(a) Mario environment (b) Atari Enduro environment
Value Meaning Value Meaning
0 No operation 0 No operation
1 Fire 1 Move up
2 Move up 2 Move right
3 Move right 3 Move left
4 Move left 4 Move down
5 Move down 5 Move up-right
6 Move up-right 6 Move up-left
7 Move up-left 7 Move down-right
8 Move down-right 8 Move down-left
9 Move down-left
10 Fire and move up
11 Fire and move right
12 Fire and move left
13 Fire and move down
14 Fire and move up-right
15 Fire and move up-left
16 Fire and move down-right
17 Fire and move down-left
(c) Atari Seaquest environment (d) Atari MsPacman environment
Table 4: Action meanings for SMB (Nguyen 2020), Atari Enduro, Seaquest and MsPacman (Brockman et al. 2016)
Appendix C Reference values used across different environments
In our experiment, we found that the PPO agents we adopted from Stable Baselines3 (Raffin et al. 2021) and Mario-PPO (Nguyen 2020) are sensitive to reference values. We tried using a constant value (zero) and blur as reference values but did not obtain clear and reasonable importance masks. However, using reference values that are visually similar to the environment background gave us better results. In Figure 4, we list the reference values we used in each environment. Specifically, for the SMB and Atari Enduro environments, we use the background color of the content in each state as the reference value. Note that the reference value for different states may vary, as the background content of each state can be different. For the Atari Seaquest environment, since all the states share the same background, we use a single reference value for each state, which combines the empty sea background and the oxygen bar background. For the Ms. Pacman environment, we simply used a constant value of the background color blue as the reference value.
Figure 4: Reference values for four environment evaluated. The content column represents a sub-figure of the original state that contains the objects and background. The last column displays the corresponding reference value for each environment.
Env Original state Content Reference value
SMB [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Enduro [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Seaquest [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
MsPacman [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Appendix D Algorithms for insertion and deletion metrics
Algorithms 1-2 present the pseudocode for evaluating the insertion and deletion metrics of (Petsiuk, Das, and Saenko 2018).
Algorithm 1 Evaluate insertion metric
1:Input: target mask $m$, state $s$, action $a$, expert policy $\pi_{E}$, fraction of pixels inserted $\alpha$, Number of pixels $N$
2:Initialize $m^{\prime}\leftarrow\text{zeros}(m)$, $p\leftarrow\emptyset$
3:$\mathcal{I}\leftarrow\text{indices}(\text{sort}(\text{flatten}(m)))$
4:for $k\leftarrow 0$ to $\lfloor\alpha N\rfloor-1$ do
5:     $(x,y)\leftarrow\mathcal{I}[k]$, $m^{\prime}[x,y]\leftarrow 1.0$
6:     $p\leftarrow p\cup\pi_{E}(s\odot m^{\prime}+r\odot(1-m^{\prime}))$
7:end for
8:return $\text{AUC}(p)$
Algorithm 2 Evaluate deletion metric
1:Input: target mask $m$, state $s$, action $a$, expert policy $\pi_{E}$, fraction of pixels deleted $\alpha$, Number of pixels $N$
2:Initialize $m^{\prime}\leftarrow\text{ones}(m)$, $p\leftarrow\emptyset$
3:$\mathcal{I}\leftarrow\text{indices}(\text{sort}(\text{flatten}(m)))$
4:for $k\leftarrow 0$ to $\lfloor\alpha N\rfloor-1$ do
5:     $(x,y)\leftarrow\mathcal{I}[k]$, $m^{\prime}[x,y]\leftarrow 0.0$
6:     $p\leftarrow p\cup\pi_{E}(s\odot m^{\prime}+r\odot(1-m^{\prime}))$
7:end for
8:return $\text{AUC}(p)$
Appendix E Object importance percentage calculation
[…]
Appendix F Time efficiency
To assess the time efficiency of mask generation, we compare the total time taken for mask generation using VisionMask against that of two baseline methods on testing datasets from four environments. Note that the time we measure does not include the training time for VisionMask. We can see that LIME and RISE are slower than VisionMask. All the experiments are conducted on A40 GPU.
Env VisionMask LIME RISE
SMB 00:00:10 25:45:02 00:51:41
Enduro 00:00:10 1:05:19 00:08:14
Seaquest 00:00:12 1:34:49 00:10:58
MsPacman 00:00:10 3:02:28 00:08:59
Table 5: Time taken to generate importance masks for three methods across four environments. The time format is hh:mm:ss. The testing dataset for SMB contains 970 states, while Atari Enduro, Seaquest, and Ms. Pacman contain 4000 states each.
Appendix G Experimental details
Dataset details
Recall from Section 3.1 that we collected the expert dataset $\pi_{E}$ with the agent being explained as expert in four envrionments. In the SMB environment(Kauten 2018), $\pi_{E}^{\text{Mario}}$ comprises a total of 9,692 state-action pairs. In Atari Enduro, Seaquest and MsPacman (Brockman et al. 2016) environment, $\pi_{E}^{\text{Atari}}$ includes 10,000 total state-action pairs. We utilize a 80%-10%-10% train-valid-test split on all expert dataset $\pi_{E}$.
Pre-processing data
We transform the RGB input to greyscale with the following function then resize it to $84\times 84$:
1def rgb_to_grayscale(img, num_output_channels: int = 1):
2 if img.ndim < 3:
3 raise TypeError(f"Input image tensor should have at least 3 dimensions, but found {img.ndim}")
4 # _assert_channels(img, [1, 3])
5
6 if num_output_channels not in (1, 3):
7 raise ValueError("num_output_channels should be either 1 or 3")
8
9 if img.shape[-3] == 3:
10 r, g, b = img.unbind(dim=-3)
11 # This implementation closely follows the TF one:
12 # https://github.com/tensorflow/tensorflow/blob/v2.3.0/tensorflow/python/ops/image_ops_impl.py#L2105-L2138
13 l_img = (0.299 * r + 0.587 * g + 0.114 * b).to(img.dtype)
14 l_img = l_img.unsqueeze(dim=-3)
15 else:
16 l_img = img.clone()
17
18 if num_output_channels == 3:
19 return l_img.expand(img.shape)
20
21 return l_img
Model details
We utilize the segmentation model DeeplabV3 as the $f_{\theta}$ model in Section 3. See the Github repo of this model from  [^1] https://github.com/pytorch/pytorch. As mentioned in the Experiment section 4, we adopt the pretrained PPO model from Stable baselines 3 (Raffin et al. 2021) as the agent to be explained. Check the Hugging Face repositories for detailed information on the PPO agents for Enduro  [^2] https://huggingface.co/sb3/ppo-EnduroNoFrameskip-v4, Seaquest  [^3] https://huggingface.co/sb3/ppo-SeaquestNoFrameskip-v4, and MsPacman  [^4] https://huggingface.co/sb3/ppo-MsPacmanNoFrameskip-v4.
Training details
We conduct the training process on a server with an A40 GPU, using 50 epochs for SMB, 36 epochs for Enduro, 52 epochs for Seaquest, and 136 epochs for Ms. Pac-Man. The optimization is done using Adam. See Table 6 for hyperparameters.
Hyperparameter value
Learning rate 0.00001
Training batch size 16
Testing batch size 1
Validating batch size 16
$\lambda_{avg}$ 0.3
$\lambda_{e}$ 1.0
$\lambda_{\text{smooth}}$ 1.0
$\lambda_{L2}$ 0.01
Table 6: Hyperparameters
Three random seeds used for training: 42, 13, 62.
Appendix H Code
Code can be found at: https://anonymous.4open.science/r/I6eV3cksE5AgZ7-EB88
Appendix I Counterfactual explanations of Super Mario Bros environment
In this section, we show more counterfactual explanations on the Super Mario Bros(Kauten 2018) environment. In Table 7 below, each row represents a random state from the environment. The original state column shows the state, the VisionMask column displays the importance map of the target action (label), and the Counterfactuals column presents the corresponding counterfactuals constructed by removing the most important regions indicated by the importance map. We detail the counterfactual explanations in Table 7 below:
  • In the first row, Mario encounters the monster Goomba and two pipes, with the bush and Goomba receiving an importance score of 31.62%, while the right pipe receives 29.2%. Counterfactual scenarios involving the removal of the Goomba and the top section of the second green pipe both result in a change of action from accelerating while jumping right to simply jumping, which may indicate that those regions are the cause of action 4 (jump right with acceleration).
  • In the second row, Mario is positioned between two brown bricks, with the left brick receiving an importance score of 25.23%, the right brick receiving 29.15%, and Mario and the bush each receiving 25.23%. It’s worth noting that jumping is necessary in the original state for Mario to pass these obstacles, but it becomes unnecessary in the counterfactual scenario. This observation suggests that the bricks may be the cause of Action 4, as when they are removed, jumping is no longer necessary.
  • The third row is identical to the example presented in the counterfactual explanation section of the Super Mario Bros (SMB) environment in the paper.
  • In the fourth column, Mario encounters a brick wall that only requires a jump (without acceleration) to pass through, with the brick wall receiving an importance score of 28.78%. The first counterfactual, removing the brick, indicates a change in action from Action 2 (right jump) to Action 4. Knowing that a right jump with acceleration (Action 4) is necessary to safely jump over the gap and land on the next platform, one can easily identify the brick wall as the cause of the action. In addition, the observation that the second counterfactual of removing the ground behind Mario did not change the action, which in turn supports our previous explanation.
Note that all the object importance percentages are calculated following E.
Table 7: Counterfactual explanations of SMB environment.
Original state VisionMask Counterfactuals
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Action 4 Action 5 Action 5
Jump right with acceleration Jump Jump
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Action 4 Action 3
Jump right with acceleration Move right with acceleration
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Action 6 Action 4 Action 1
Move left Jump right with acceleration Move right
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Action 2 Action 4 Action 2
Jump right Jump right with acceleration Jump right
Appendix J Counterfactual explanations of Atari Enduro environment
In this section, we presents counterfactual explanations on the Atari Enduro environment. In Table 8 below, each row represents a random state from the environment. The original state column shows the state, the VisionMask column displays the importance map of the target action (label), and the Counterfactuals column presents the corresponding counterfactuals constructed by removing the most important regions indicated by the importance map. We detailed the counterfactual explanations in Table 8 below:
  • In the first row, the agent attempts to move left to avoid a collision with the incoming blue car, which receives an importance score of 8.33% as indicated by the importance map of second column. The counterfactual scenario involving the removal of the blue car leads to a change in action from moving left to accelerating while moving left. It’s important to note that accelerating while moving left could still result in a collision with the blue car in the original state.
  • In the second row, the agent attempts to overtake the blue car on the right, which receives an importance score of 11.25%. The corresponding counterfactual scenario of removing the blue car indicates a transition of action from accelerating to moving right and accelerating. Given that moving right with accelerating will cause a collision with that car, we may attribute the accelerating action to the blue car.
  • In the third row, a collision occurs in the original state with Action 3 (moving left). The agent tries to move left to avoid the collision but fails, where the collided blue car receives an importance score of 16.10%. By removing the blue (collided) car from the original state, the action in that counterfactual scenario changes to moving left with accelerating One may explain the collided blue car as the main cause of the action (moving left) in this counterfactual scenario.
  • In the last row, the agent tries to turn right, where the race track receives an importance score of 27.58%. In the counterfacutal scenario, the action changes from moving right with acceleration to moving left with acceleration when most of the race track is removed from original state. It’s obvious that the right directioned race track is the main cause of action of truing right with accelerating.
Table 8: Counterfactual explanations of Atari Enduro environment. Note that the collision happens in the original state in the third row.
Original state VisionMask Counterfactuals
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Action 3: left Action 8: left + accelerate
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Action 1: accelerate Action 7: right + accelerate
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Action 3: left Action 8: left + accelerate
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Action 7: right + accelerate Action 8: left + accelerate
Appendix K Counterfactual explanations of Atari Seaquest environment
In this section, we presents counterfactual explanations on the Atari Seaquest environment. In Table 9 below, each row represents a random state from the environment. The original state column shows the state, the VisionMask column displays the importance map of the target action (label), and the Counterfactuals column presents the corresponding counterfactuals constructed by removing the most important regions indicated by the importance map. We detailed the counterfactual explanations in Table 9 below:
  • In the first row, the agent attempts to attack the shark in the bottom left corner, marked in green, while the shark receives an importance score of 5.65%. In the counterfactual scenario where the shark is removed from the original state, the resulting action shifts from moving down and firing to moving down and to the right. The absence of the shark reasonably leads to a shift in action from firing to moving right, indicating that the presence of the shark likely prompted the original action.
  • The second row is similar to first row where the counterfactual of removing shark could incur the action shifting from fire to a movement.
  • In the last row, the agent’s submarine is about to start diving with a full tank of oxygen, where the oxygen bar receives most of the importance score, 53.51%. Counterfactual of removing the half of the oxygen bar making the action change from moving down, left and fire to moving left to stay on the surface. The oxygen level can easily be explained as the main cause in that counterfactual scenario.
Table 9: Counterfactual explanations of Atari Seaquest environment
Original state VisionMask Counterfactuals
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Action 13: down + fire Action 8: down + right
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Action 10: up + fire Action 6: up + right
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Action 17: down + left + fire Action 4: left